OpenStudy (anonymous):

solve for m: 2/3 = m/42 I got m=28 but someone eles said m=16

OpenStudy (anonymous):

\[\frac{ 2 }{3 }=\frac{ m }{ 42 }\] \[\frac{ 2 }{ 3 }\times42=m\] \[28=m\] Yup, your answer is correct! m=28

OpenStudy (anonymous):

ok thanks